# VURF: A General-purpose Reasoning and Self-refinement Framework for Video Understanding

### v2

@@ 4.1 Video Question Answering
VQA is an important task in video comprehension that endeavors to connect natural language processing with video comprehension. The objective is to empower models to interpret the content of a video and respond to user queries related to the video content. Effective resolution of such queries often involves breaking down the overarching problem into manageable subproblems.
For instance, consider a video of a man entering a room and then performing some action. If a human is asked, “What does the man do after entering the room?” they would visually inspect the video, locate the man, observe his actions by examining the video and then deduce what the nature of his activities. This example illustrates the critical role of logical reasoning in decomposing complex questions into more manageable sub-queries, suitable for evaluation using off-the-shelf models. The inherent complexity of such tasks makes VQA an ideal candidate for a visual programming approach, given its primary function of decomposing intricate tasks into more manageable components. In the context of the aforementioned example, VURF, as demonstrated in Fig. 7 (Right), initially employs a GROUNDING model to identify an interval where the man enters the room. Subsequently, it would call the TRIMAFTER module to retrieve the relevant part of the video and finally using VQA module, it will engage in visual examination of the video to answer the question “Pick up towel”. This approach makes the task more logical and interpretable, with possible explanations in case an output is wrong.
VURF also extends its functionality to Multi-Video Question Answering tasks which involves synthesizing information from two distinct videos to provide accurate responses to user queries. Our approach simplifies this intricate process by breaking it down into a sequence of steps executed by a language model. For instance, in Fig. 5, when presented with the user query “Are the number of people in the kitchen in both videos the same?”, our approach generates a visual program. This program initiates the TRACK step in both videos to identify and isolate regions containing people. Subsequently, the COUNT step computes the number of people in each video, and a final answer is derived through a comparison of these counts.
Figure 6: Qualitative example showing the program steps of the Pose Estimation task of VURF. We show limited frames due to space constraints. Best viewed in zoom
@@ 4.2 Pose Estimation
Pose Estimation identifies the position and orientation of human bodies within images or videos and uses a skeletal model to represent them [78]. This field has drawn significant attention in the research and has been utilized in various domains such as HCI [16], virtual reality [45], clinical use in pediatric disorders [17], and motor assessment in adults [53]. Numerous methods have been employed to solve pose estimation. The general workflow mainly operates in two stages, first, identify and locate the orientation of joints in the human body e.g., by using ‘Human model’ in MMPose [50], neural network architecture in OpenPose [8], and through the 3D-2D Keypoints Matching [10]. The next step is to use a framework or model that inputs this graph skeleton and outputs the estimated pose of the object [3].
Although these model-based approaches in pose estimation frameworks perform well on numerous problems, the aforementioned approaches can not be generalized and they fail to generate reproducible results on unseen datasets. The application of pose estimation becomes a natural choice in our Visual Programming approach as it essentially pre-processes the video, focuses on specific human being (using the TRACK and CROP functionality) and then performs the pose classification. In our framework, we utilize the ‘Human model’ from MMPose [50] for keypoint detection in video frames. For example, as shown in Fig. 7 (Left) which depicts the person’s pose tagging in the video.
We then use these key points to classify the pose of a person (e.g. if a person is falling). This concept can be extended to hazard detection, street-crime detection, fight detection, yoga pose detection, and other applications.
For instance, consider a scenario in a video feed where a person is falling at the end Fig. 6. The user question, "What happens to the man at the end of the video?" will be decomposed into sub-tasks. It starts with triggering the TRIM step to filter out the relevant video frames only. This is followed by the POSE DETECT function applied to the preceding output frames from step 1. The output of POSE DETECT then acts as input to the succeeding CLASSIFY step, which detects that a man is falling.
@@ 4.3 Video Editing
Video editing is a crucial process in the post-production phase of film-making, television production, and other visual media industries [14]. It involves manipulating and rearranging video clips to create a coherent and engaging narrative or visual presentation. Video editing encompasses various tasks such as trimming, cutting, merging, and arranging video segments, as well as adding visual effects, transitions, and audio enhancements. In today’s digital age, video editing is typically performed using specialized software applications that offer a wide range of tools and features to facilitate the editing process.
@@
Figure 7: Qualitative examples demonstrating three different use cases of the VURF framework including video editing, pose estimation and video question answering. The reasoning framework effectively understands the instructions and breaks down the problem into sub-tasks that are solved to reach the final outcome. Left: An example of Pose Detection that uses the visual program generated by VURF. Middle: A qualitative example of Video Editing which directly leverages VURF’s list of modules. Right: An example of VQA task decomposition by manageable sub-tasks by our framework. Best viewed in zoom.
@@ 5 Experiments and Results
Our experiments encompass both quantitative and qualitative evaluation of the proposed video reasoning framework. In the quantitative experiments, we assess the performance impact of the proposed video programming approach built on a pretrained model designed specifically for the Video Question Answering task. Additionally, we examine how the self-refinement approach affects the performance of the video programming approach, exploring the effects of altering the number of iterations on the model outputs. Finally, we present qualitative examples involving diverse tasks such as Pose Detection and Video Editing to highlight the efficacy of video programming in the context of video understanding tasks.
@@ 5.1 VQA Evaluation
In recent research endeavors, diverse approaches to video question answering (VQA) have been investigated, yielding promising outcomes through the utilization of large pre-trained vision-language models. Notably, the SeViLA framework [72], has emerged as a distinguished approach, effectively addressing important challenges in video processing. SeViLA integrates temporal keyframe localization and question answering by employing a unified image-language model (BLIP2)[35]. Demonstrating strong performance on VQA task, SeViLA achieves state-of-the-art accuracy across various VQA datasets, including STAR [63], NextQA [65], and Social IQ QA [74]. This paper incorporates SeViLA as the VQA module, showcasing improved performance in the zero-shot video question answering task with our VURF. The assessment is conducted on four benchmark datasets: STAR, NextQA, Social-IQ QA, and TrafficQA [66].
@@
Several approaches, including InternVideo [62] and ViperGPT [55], have been proposed to address the challenge of Visual Question Answering alongside SeViLA. However, as illustrated in Table 1, VURF outperforms all three in zero-shot video question answering. Leveraging SeViLA as our VQA module provides a natural baseline for comparison. VURF offers a significant advantage over SeViLA and InternVideo by enabling reasoning through instructions rather than functioning as a black box, facilitating self-improvement. This capability proves advantageous over ViperGPT, which, despite employing the reasoning power of LLMs, is outperformed by VURF. The primary reason for this discrepancy is ViperGPT’s susceptibility to contextual hallucinations. We demonstrate that integrating self-refinement of in-context program-instruction pairs and incorporating a simple error correction module substantially enhances the performance of a visual programming approach.
While the concept of breaking down instructions or questions into simpler steps is generally beneficial, its effectiveness heavily relies on the functionality of the utilized functions. This limitation is evident in NextQA [65], particularly in categories involving temporal relations. Initially, we employed the GROUND functionality utilizing the TubeDETR [67] model. However, our observations revealed that when utilizing the GROUND/Localisation function, VURF does not surpass the baseline SeViLA model, primarily due to inaccuracies in the temporal localization functionality. To address this limitation and showcase the efficacy of the video programming approach, we manually annotated ground truths for the localization functionality in 30 randomly selected videos, resulting in improved accuracy (see Table 2). This underscores that with precise Temporal Localization functionality, VURF surpasses the performance of the base model. For the conducted testing, as depicted in Table 1, we deliberately refrain from utilizing the GROUND functionality to ensure fairness in the assessment. Although this omission slightly decreases the accuracy from 65.2% to 64.0%, it still exceeds that of SeViLA.
Finally, since the Social-IQ-2.0 dataset [74] contains the transcripts of the videos so we introduce a new function, ANALYSE, tailored for questions aimed at discerning the mood or tone of the video conversation. This function involves querying a Language Model (LLM) with the transcript, and, in some instances, the appended summary along with the posed question. Our analysis includes performance metrics for both scenarios i.e., employing transcripts and not using them (Table 3).
@@ 5.2 VQA Ablations
ViperGPT [55] diverges from VURF by not incorporating self-refinement mechanisms. This distinction is noteworthy as it highlights differing approaches to handling instruction comprehension within the context of VQA. Sole reliance on pre-trained language models without iterative refinement mechanisms makes the approach susceptible to contextual hallucinations. These hallucinations, stemming from the model’s inability to accurately contextualize instructions within the visual domain, can lead to erroneous outputs and decreased performance in tasks such as zero-shot video question answering. Our ablations (see Table 4) clearly demonstrate that integrating the self-refinement of in-context program-instruction pairs and incorporating a simple error correction module substantially enhances the performance of a visual programming approach.
Table 1: Performance of VURF on Video Question Answering (zero-shot) compared to other existing models. Each result is evaluated on the validation set of the corresponding dataset. The table shows the accuracies (in percentage) of the respective experiments.
@@
InternVideo [62] 50.2% 41.8% 30.1% 31.2%
ViperGPT [55] 60.0% 40.3% 37.8% 35.7%
SeViLA [72] 63.8% 44.3% 47.3% 39.1%
@@
Table 2: Demostration of VURF without the influence of an imperfect module. We evaluate the performance of our Video Question Answering task on 30 randomly selected data points from the NextQA dataset by first using an imperfect temporal localization module (GROUND) and then replacing it by the actual ground truths of the localization queries. We find a noticeable increase in performance once a perfect Temporal Grounding model is introduced.
Dataset No Ground Truth Ground Truth
NextQA [65] 61.4% 65.2%
Table 3: Effect of using transcripts from the Social-IQ QA dataset. We evaluate the performance of our Video Question Answering task on Social-IQ QA dataset. The dataset contains transcripts so we conduct two experiments; with and without using transcripts. The results show an increase in performance (as expected) when the transcripts are utilized.
Dataset Baseline With transcripts
Social-IQ-2.0 [74] 48.1% 51.6%
@@ 5.3 Self-Refinement
    In-Context Example Refinement: In examining the self-refinement process, we employ the NeXt-QA dataset by randomly sampling 50 videos for the test set. Additionally, we curate a set of 20 in-context examples for evaluation. The accuracy of the test set is initially calculated. Subsequently, these in-context examples undergo self-refinement. This approach initially presents instructions to an LLM for program generation without contextual influence. The generated program, along with the original program, is then provided to the LLM with a prompt to maintain the structure of the initial program while enhancing it using the non-contextual program. The refined in-context examples undergo multiple iterations of this process, and we report the accuracy on the test set with respect to the number of iterations of self-refinement applied to the in-context examples in Fig. 7(a).
@@
    (a) Accuracy plotted against a number of iterations of self-refinement We iteratively refine our in-context examples for a 50-video sample of NextQA and report our accuracy against the number of iterations of this self-refinement. The accuracy appears to quickly stabilize after 1 or 2 iterations.
    (b) Number of errors plotted against a number of iterations of self-refinement. The plot shows how the number of errors in the generated programs of a sample of 400 video-instruction pairs from the STAR dataset is affected by the number of iterations of the auto-error correction module.
@@ 6 Conclusion
Appendix
7 Dataset Specific Prompts
We curated in-context examples for 4 different datasets (NeXTQA, STAR, Social-IQ, and TrafficQA). The prompts for two of the datasets are shown in Figure 9. We show 3 in-context examples for each prompt.
Figure 9: LEFT: LLM prompt for STAR dataset RIGHT: LLM prompt for TrafficQA dataset
8 Refinement Prompts
8.1 Error Correction
The error correction module queries the LLM two times, one to receive feedback on a program and the other to correct the program given the feedback. The prompts are shown in Figure 10. We display 2 in-context examples for each prompt.
Figure 10: LEFT: The prompt for the feedback generation of a given program. RIGHT: Given a feedback the correct program is generated by giving this prompt to the LLM.
8.2 Self-Refinement
The self-refinement module (which we pre-apply to refine our in-context examples) consists of 2 major queries to the LLM. One is for generating a context-free program to avoid hallucinations. Other is to convert the context-free program to a valid program. The prompts are shown in Figure 11
Figure 11: TOP: The prompt for the context-free generation. BOTTOM: The prompt for aligning a "perfect" program to a valid program
Figure 12: An example of the self refinement module. An initially generated prompt is refined by the LLM to produce a more complex and modular program.
Figure 13: A qualitative example of video-editing using VURF.
Figure 14: A step by step qualitative example of Video Question Answering using VURF.

### v3

@@ 4.1 Video Question Answering
VQA is an important task in video comprehension that endeavors to connect natural language processing with video comprehension. The objective is to empower models to interpret the content of a video and respond to user queries related to the video content. Effective resolution of such queries often involves breaking down the overarching problem into manageable sub-problems.
For instance, consider a video of a man entering a room and then performing some action. If a human is asked, “What does the man do after entering the room?” they would visually inspect the video, locate the man, observe his actions by examining the video, and then deduce what the nature of his activities. This example illustrates the critical role of logical reasoning in decomposing complex questions into more manageable sub-queries, suitable for evaluation using off-the-shelf models. The inherent complexity of such tasks makes VQA an ideal candidate for a visual programming approach, given its primary function of decomposing intricate tasks into more manageable components. In the context of the aforementioned example, VURF, as demonstrated in Fig. 7 (Right), initially employs a GROUNDING model to identify an interval where the man enters the room. Subsequently, it would call the TRIMAFTER module to retrieve the relevant part of the video, and finally using the VQA module, it will engage in visual examination of the video to answer the question “Pick up towel”. This approach makes the task more logical and interpretable, with possible explanations in case an output is wrong.
VURF also extends its functionality to Multi-Video Question Answering tasks which involves synthesizing information from two distinct videos to provide accurate responses to user queries. Our approach simplifies this intricate process by breaking it down into a sequence of steps executed by a language model. An example is shown in Fig. 5.
Figure 6: Qualitative example showing the program steps of the Pose Estimation task of VURF.
@@ 4.2 Pose Estimation
Pose Estimation identifies the position and orientation of human bodies in images or videos using a skeletal model and has been applied in areas like HCI, virtual reality, and clinical assessments Zheng et al. (2023); Erol et al. (2007); Escobar et al. (2019). The process typically involves two stages: detecting joint orientations, as seen in tools like MMPose Sengupta et al. (2020) and OpenPose Cao et al. (2017), followed by using a model to estimate the pose from the skeletal representation Andriluka et al. (2014). While these approaches are effective, they often struggle to generalize to unseen datasets.
In our Visual Programming approach, pose estimation is used to pre-process videos by tracking and cropping specific individuals, followed by pose classification. We utilize MMPose for keypoint detection and classify poses for tasks like fall detection, which can be extended to applications such as hazard or crime detection. For example, Fig. 7 shows pose tagging in a video, and Fig. 6 illustrates a fall detection scenario where relevant frames are trimmed, pose is detected, and the fall is classified.
@@ 4.3 Video Editing
Video editing is a crucial process in the post-production phase of film-making, television production, and other visual media industries Dancyger (2018). It involves manipulating and rearranging video clips to create a coherent and engaging narrative or visual presentation. Video editing encompasses various tasks such as trimming, cutting, merging, and arranging video segments, as well as adding visual effects, transitions, and audio enhancements. In today’s digital age, video editing is typically performed using specialized software applications that offer a wide range of tools and features to facilitate the editing process.
@@
Figure 7: Qualitative examples demonstrating different use cases of VURF. Left: An example of Pose Detection that uses the visual program generated by VURF. Middle: A qualitative example of Video Editing that directly leverages VURF’s list of modules. Right: An example of VQA task decomposition by manageable sub-tasks by our framework. Best viewed in zoom.
@@ 5 Experiments and Results
Our experiments encompass both quantitative and qualitative evaluation of the proposed video reasoning framework. In the quantitative experiments, we assess the performance impact of the proposed video programming approach built on a pre-trained model designed specifically for the Video Question Answering task. Additionally, we examine how the self-refinement approach affects the performance of the video programming approach, exploring the effects of altering the number of iterations on the model outputs. Finally, we present qualitative examples involving diverse tasks such as Pose Detection and Video Editing to highlight the efficacy of video programming in the context of video understanding tasks.
@@ 5.1 VQA Evaluation
Diverse approaches to video question answering (VQA) have been investigated, yielding promising outcomes. However, the SeViLA framework Yu et al. (2024), has emerged as a distinguished approach, integrating temporal keyframe localization and question answering by employing a unified image-language model (BLIP2)Li et al. (2023). We incorporate SeViLA as the VQA module, showcasing improved performance in the zero-shot video question-answering task with our VURF. The assessment is conducted on four benchmark datasets: STAR, NextQA, Social-IQ QA, and TrafficQA Xu et al. (2021).
@@
Several approaches, including InternVideo Wang et al. (2022) and ViperGPT Surís et al. (2023), have been proposed to address the challenge of Visual Question Answering alongside SeViLA. However, as illustrated in Table 1, VURF outperforms all three in zero-shot video question answering. VURF offers a significant advantage over SeViLA and InternVideo by enabling reasoning through instructions rather than functioning as a black box, facilitating self-improvement. This capability proves advantageous over ViperGPT, which, despite employing the reasoning power of LLMs, is outperformed by VURF. The primary reason for this discrepancy is ViperGPT’s susceptibility to contextual hallucinations. We demonstrate that integrating the self-refinement of in-context program-instruction pairs and incorporating a simple error correction module substantially enhances the performance of a visual programming approach.
@@ 5.2 VQA Ablations
ViperGPT Surís et al. (2023) diverges from VURF by not incorporating self-refinement mechanisms. This distinction is noteworthy as it highlights differing approaches to handling instruction comprehension within the context of VQA. Sole reliance on pre-trained language models without iterative refinement mechanisms makes the approach susceptible to contextual hallucinations. Our ablations (see Table 3) clearly demonstrate that integrating the self-refinement of in-context program-instruction pairs and incorporating a simple error correction module substantially enhances the performance of a visual programming approach.
Moreover, we also compare our method with visual instruction tuning Liu et al. (2023). To this end, we finetune a LlaVA-based model (trained using Visual instruction tuning) called Video-ChatGPT Maaz et al. (2024) on question-program pairs and test the performance of generated programs on the STAR VQA dataset. VURF outperforms the visual instruction tuning method (Table 2), even when the latter has access to more data during the training phase compared to the number of in-context examples used by VURF. VURF’s superior performance stems from its dynamic adaptation to various scenarios, leveraging contextual cues to manage tasks on the fly.
Table 1: Performance of VURF on Video Question Answering (zero-shot) compared to other existing models. Each result is evaluated on the validation set of the corresponding dataset.
@@
InternVideo Wang et al. (2022) 50.2% 41.8% 30.1% 31.2%
ViperGPT Surís et al. (2023) 60.0% 40.3% 37.8% 35.7%
SeViLA Yu et al. (2024) 63.8% 44.3% 47.3% 39.1%
@@
Table 2: Comparison with Visual Instruction Tuning. We finetune Video-ChatGPT on program-question pairs (which we use as in-context examples in our system) and then test their program-generating performance.
Dataset VURF (ours) Visual Instruction Tuning
STAR $44.5\%$ $22.3\%$
@@ 5.3 Self-Refinement
    In-Context Example Refinement: In examining the self-refinement process, we employ the NeXt-QA dataset by randomly sampling 50 videos for the test set. Additionally, we curate a set of 20 in-context examples for evaluation. The accuracy of the test set is initially calculated. Subsequently, these in-context examples undergo self-refinement. This approach initially presents instructions to an LLM for program generation without contextual influence. The generated program, along with the original program, is then provided to the LLM with a prompt to maintain the structure of the initial program while enhancing it using the non-contextual program. The refined in-context examples undergo multiple iterations of this process, and we report the accuracy on the test set concerning the number of iterations of self-refinement applied to the in-context examples in Fig. 8(a).
@@
    (a) Accuracy plotted against the number of iterations of self-refinement.
    (b) Number of errors plotted against the number of iterations of self-refinement.